OpenStudy (anonymous):

300 cm3 of a 2.00 m solution kno3 is mixed with 200 cm3 of a 3.00 solution of kno3. whats the concentration of the resulting solution

OpenStudy (matt101):

We need to find the total moles of KNO3 first, which we will then use to find the concentration of the resulting solution. A cm^3 is a mL, so 300 cm^3 = 300 mL = 0.3 L. 0.3 x 2 = 0.6 mol in the first solution. Similarly, 0.2 x 3 = 0.6 mol in the second solution. 0.6 + 0.6 = 1.2 total moles of KNO3. The total volume is 0.2 + 0.3 = 0.5 L. This means the final concentration is 1.2/0.5 = 2.4 M.

OpenStudy (anonymous):

thanks